show that linear combination of eikx and e-ikx is a eigen function of d2/dx2

1
Expert's answer
2021-10-05T10:07:41-0400

Let us consider the linear combination "\psi(x) = C_1 e^{i k x} + C_2 e^{- i k x}", where "C_1, C_2" - constants. Then,

"\frac{d^2}{d x^2} \psi(x) = C_1 (i k)^2 e^{i k x} + C_2 (-i k)^2 e^{-i k x} = -k^2(C_1 e^{i k x} + C_2 e^{-i k x}) = - k^2 \psi(x)".

Hence, "\psi(x)" is an eigenfunction of differential operator "\frac{d^2}{d x^2}" with corresponding eigenvalue "-k^2".